If a crew of 8 can row a boat 20k/hr in still water. The river flows 2K/hr. If they row the boat up stream and return in 15min.how far up stream did they go?
Answer by ptaylor(1894)

Let d=distance they rowed upstream
distance they rowed upstream = distance they rowed downstream
speed (rate) upstream equals speed of boat in still water (20k/hr) minus speed of current(2k/hr)=18k/hr
speed (rate) downstream equals speed of boat in still water (20k/hr) plus speed of current(2k/hr)=22k/hr
distance(d)=rate(r) times time (t) or d=rt and t=d/r
time upstream =d/18
time downstream=d/22
Now we are told that time upstream (d/18) plus time downstream (d/22)=15 minutes or 0.25 hr. so our equation to solve is
d/18+d/22=0.25 Multiply each term by (18)(22) to get rid of fractions
22d(18/18)+18d(22/22)=0.25(18)(22) clear parens
22d+18d=99
40d=99
d=2.475 km---------------distance upstream
ck
(2.475)/18+(2.475)/22=0.25
0.1375+0.1125=0.25
0.25=0.25

Hope this helps----ptaylor